# Kac–Moody Symmetric Spaces: arbitary symmetrizable complex or almost split real type

### v3

@@ 4.4 Flats
Let $G$ be a split Kac–Moody group over $\mathbb{K}$ of type $\mathbb{A}$ and equip the real form $\mathfrak{a}$ with its Euclidean symmetric space structure (cp. 2.2 1). Then the map
@@
Now, using that the action of $G$ on the symmetric space $G/K$ is given by an automorphism (cf. 4.4), we can assume that $g=e$. Let $X,Y\in\mathfrak{a}$ and use that for all $t\in A$ follows that $\Theta(t)=t^{-1}$ (cp. Section 3.5). Then one computes
@@
In claim $5$ a reductive split real Lie group is used and the fact that it is a subgroup of $\operatorname{GL}_{m}(\mathbb{R})$. For $\mathbb{K}=\mathbb{C}$ we also have an algebraic group which is a subgroup of $\operatorname{GL}_{m}(\mathbb{C})$ (cp. [Bor91, Proposition 1.10]. As we have seen in the proof of 4.1, the involution $\Theta$ at the level of the matrix group can be understood as transpose, inversion, which is combined with complex conjugation (cp. [PR94, Theorem 3.7 and Proposition 3.9]). In the concrete calculation of the product $xy\cdot xy$, the upper left entry is equal to a sum of the squares of the absolute values of the entries in the first column. Now the arguments in claim $5$ work verbatim exactly as in the real case. ∎
@@ 4.4.1 Flats in the Group Model
Recall from Section 4.3 the isomorphism $\hat{\tau}\colon G/K\to\tau(G)$ between the coset and the group model and that it is equivariant under the action of $G$. We are now interested in the image of a standard flat in the group model under this isomorphism:
@@
$=g\tau(A)\Theta(g)^{-1}$
$=g\star\tau(A).$
Thus, the image of the standard flat $gAK$ in the group model is given by $g\star\tau(A)$. Therefore, denote by $F[g]\coloneqq g\star A\subseteq\tau(G)$ the standard flat of the group model (cp. [FHHK20, Proposition 5.4]). In order to establish a relationship between the flats in the symmetric space $G/K$, the twin apartments in the corresponding building $\Delta_{\pm}$ and the maximal tori of $G$, we need the following lemma.
@@
Let $G$ ba a split Kac–Moody group over $\mathbb{K}$ of arbitrary type. Then
@@
Note that $N_{K}(T)<N_{K}(A)$, based on the fact that the conjugation is continuous and $A$ is the identity component of $T$.
For the reverse inclusion, consider an element normalizing $A$, then the conjugation with this element shifts $T$ to a maximal torus $T^{\prime}$ of $G$. But the standard maximal torus $T$ containing $A$ is unique ([Cap09, Lemma 4.9]). It follows that $T=T^{\prime}$, thus this element normalizes $T$. ∎
@@
First denote by $T_{max}$ the set of all maximal tori of $G$ and with $F_{std}$ the set of all standard flats in $G/K$. Then there is a $G$-equivariant map between these sets given by
@@
Together with the fact that $N_{G}(T)$ acts trivially by conjugation on $T_{max}$, it follows that $G/N_{G}(T)$ can be identified with $T_{max}$ by mapping $gN_{G}(T)\mapsto gTg^{-1}$, in particular $f$ is surjective.
@@
Note that $T$ is the stabilizer of the standard twin apartment and the group $G$ acts transitively on $T$ and the twin apartments. For the action of $G$ on $T$ the stabilizer of the standard torus is given by $N_{G}(T)$ and for the action of $G$ on the twin apartments is $N_{G}(T)$ the stabilizer of the standard twin apartment. Thus, for the twin apartments follows that the maximal tori of $G$ are exactly the chamberwise stabilizers of the twin apartments of $\Delta_{\pm}$. ∎
@@ 5 Automorphism of Kac–Moody Symmetric Spaces
As a next step in developing the theory of symmetric spaces for Kac–Moody groups over $\mathbb{K}$ of type $\mathbb{A}$, one can analyzes the automorphism group of the whole symmetric space and the local action on flats. For the non-affine real case, in chapter 6 of [FHHK20], the authors observe several statements for the global automorphism group on the symmetric space as well as for the local automorphism group with respect to the flats.
@@ 5.1 Global Action
We are going to use [FHHK20, Proposition 2.33]: If a subgroup of the automorphism group of the symmetric space acts, which has to contain the transvection group, transitively on the maximal flats, then the group acts strongly transitively. This statement is shown for abstract symmetric spaces, hence we can apply it here. In 4.7 we have seen that $G$ contains the transvection group, so we can use the statement quoted.
By 4.12 we already know that $G$ acts transitively on standard flats and by 4.10 all standard flats are maximal flats. Therefore, the statement follows. ∎
A central assumption in [FHHK20] which allows to decompose an automorphism, is that the Kac–Moody group is centered (cp. 3.13), i.e. the Kac–Moody group is generated only by its root groups. To establish a similar results for non-centered Kac–Moody groups $G$ of type $\mathbb{A}$, we have to take care of the extended torus. In order to do this, we consider in the following section $G^{\dagger}\trianglelefteq G$.
@@
To do this, we need to recall some thins from Section 3.2. First, the map $\eta$ from the Tits functor, further the fact, that an element of the root group $U_{\alpha}$ is given by $x_{\alpha}(r)$ for $r\in\mathbb{K}\in\{\mathbb{R},\mathbb{C}\}$ and $\alpha\in\Delta^{re}$. And lastly, that the torus of a Kac–Moody group $T=\mathfrak{T}_{\Lambda}(\mathbb{K})$ is isomorphic to $\operatorname{Hom}_{grp}(\Lambda,\mathbb{K}^{\times})$. Hence we can define the map
@@
By the discussion in [CM05, Section 8.2] it follows that the sign automorphism is unique and from [Cap09, Section 8.2.1] follows that it is related to the Chevalley involution. It follows that for $\mathbb{K}=\mathbb{R}$ one has
@@
The case $k=\mathbb{R}$ is proven in [FHHK20, Proposition 6.5], where it is formulated for any generalized Cartan-Matrix $\mathbb{A}$.
For the complex case, we recall the decomposition of an automorphism of $G^{\dagger}$ ([CM05, Theorem 2.4]: Denote with $c_{g}$ an inner automorphism, with $\omega$ the Chevalley involution, with $d$ a diagonal automorphism, with $\gamma$ a diagram automorphism and with $f$ a field automorphism. Then any automorphism $\alpha\in G^{\dagger}$ can be written as
@@
In the complex case, there is only one non-trivial continuous field automorphism $f$, the complex conjugation. In particular, $f$ commutes with all other automorphism in this composition and the generated subgroups of these homomorphisms intersect trivially. By the definition of the involution $\Theta$ follows $\omega\circ f=\Theta$. ∎
@@
  • Since we are interested in statements about Kac–Moody symmetric spaces, we also need to define a symmetric space with respect to the centered subgroup $G^{\dagger}$. In particular, let $K^{\dagger}\coloneqq G^{\dagger}\cap K$, and since $\Theta$ is an automorphism of $G$, it follows that $(G^{\dagger})^{\Theta}=K^{\dagger}$. Now define the corresponding Kac–Moody symmetric space
@@
We conclude this section by pointing out the following statement, which describes the automorphism group of a Kac–Moody symmetric space coming from the centered subgroup. Since $X^{\dagger}$ corresponds to the unreduced Kac–Moody symmetric space in [FHHK20], the proof of this theorem is the same as that of the quoted theorem. Note that the underlying field plays no role in the proof of this statement.
@@ 6 Local Action
  • Call a point $q\in F$ singular with respect to $p$, if there exists a second maximal flat $E\subseteq X$ such that $p,q\in E$ and $E\neq F$. If not, call $q$ regular. Further, let ${F^{\text{sing}}}(p)$ denote the subset of singular points of $(p,F)$ and ${F^{\text{reg}}}(p)$ the set of regular points. It follows that the flat $F$ can be written as $F={F^{\text{reg}}}(p)\sqcup{F^{\text{sing}}}(p)$.
@@
From the non-degeneracy of B on $\mathfrak{h}$ it follows that B is also non-degenerate on $\mathfrak{a}$. Therefore, the reflection $r_{\alpha}^{\vee}$ under $\rho_{KM}$ is the only B-orthogonal reflection on the hyperplane $H_{\alpha}$. In particular, this leads to an one-to-one correspondence between the real roots $\alpha\in\Delta^{re}$ and the hyperplanes $H_{\alpha}$.
@@
As we saw at the beginning of this section, we can also define singular points on $\mathfrak{a}$ by using the notion of singular points in the context of abstract symmetric spaces. To be able to compare the two terms, we need to make them compatible. Therefore, let $(e,AK)$ be the standard flat of the symmetric space $G/K$. Using the chart $\varphi_{e}$, one can move the singular points of the flat $AK$ into the real form $\mathfrak{a}$, i.e.
@@
Since $G$ acts strongly transitively, see 5.1, any pointed maximal flat $(p,F)$ can mapped to the standard pointed flat $(e,AK)$. Furthermore, the $G$ action preserves the decomposition into regular and singular points. Thus we observe an isomorphism
@@
In order to be able to use some corollaries from the appendix of [FHHK20], we need an additional assumption in the rest of this section: the generalized Cartan matrix $\mathbb{A}\in\mathbb{Z}^{n\times n}$ must be symmetrizable and $n$ must be at least $2$.
@@
Note that [FHHK20, Corollary A.37] is applicable here, since the arguments and statements they used in the proof of this corollary are based on the fact that the bilinear form B coming from the generalized Cartan matrix $\mathbb{A}$ is non-degenerate and that $\rho_{KM}$ is faithful. All these properties are also satisfied in our context and thus we can apply the corollary.
@@
Therefore given a pointed flat $(p,F)$ and a local transformation $f\in\operatorname{GL}(p,F,{F^{\text{sing}}})$, then $f$ is a local automorphism if $\varphi^{-1}\circ f\circ\varphi$ is B-orthogonal for a chart $\varphi$ centered at $p$. We denote the subgroup of local automorphism by $\operatorname{Aut}(p,F)<\operatorname{GL}(p,F,{F^{\text{sing}}})$. By the previous discussion, we can decompose any local transformation into a product of a homothety and a B-orthogonal transformation. It follows that we can now decompose the group of local transformation as follows
@@
$\operatorname{Aut}(\Sigma)=W\rtimes\operatorname{Aut}(\mathcal{W},S)$
@@ 6.1 Comparing the Weyl Groups
Recall that the extended Weyl group $\widetilde{\mathcal{W}}\coloneqq\langle\tilde{s}_{i}\mid i\in I\rangle$ (cp. Section 3.2) surjects on $\mathcal{W}$.
@@
where $A^{\dagger}$ and $M^{\dagger}$ are defined as in 3.14. By the construction of $G^{\dagger}$, it follows that $T^{\dagger}\cong(\mathbb{R}^{\times})^{n}$, where $n$ is the size of the Cartan matrix $\mathbb{A}$. Hence it follows by the construction of $M^{\dagger}$ (cp. 3.14) that $M^{\dagger}$ is the unique maximal finite subgroup of elements of order $2$. This implies the following
@@
Altogether it follows that
@@
On the other hand, one knows that
$\nicefrac{{\widetilde{\mathcal{W}}M}}{{M}}\cong\nicefrac{{\widetilde{\mathcal{W}}}}{{(\widetilde{\mathcal{W}}\cap M)}}=\nicefrac{{\widetilde{\mathcal{W}}}}{{M^{\dagger}}}\cong\mathcal{W}$
and thus
@@
Now we analyze the stabilizer and fixator of $G/K$ with respect to the action of $G$ on $G/K$. Observe that the fixator of $e$ is given by $K$ and $N_{G}(T)$ is the stabilizer of $AK$ (the calculation is done in the proof of 4.12). Thus $\operatorname{Stab}_{G}(e,AK)=N_{K}(T)$ follows and by 6.4 follows that $\widetilde{\mathcal{W}}M\cong N_{K}(T)$.
By 3.14 (A3) follows that $M$ is a subset of $K$ and centralizes $A$. Hence $M$ is a subgroup of $\operatorname{Fix}_{G}(p,F)$. Using 6.4 and the isomorphism Equation 10 it follows that
@@
So far we have established the isomorphism $\operatorname{Stab}_{G}(e,AK)/M\cong\mathcal{W}$. Since the representation $\rho_{KM}\colon\mathcal{W}\to\operatorname{GL}(\mathfrak{a})$ (cp. Equation 7) is faithful, each element different from the identity of $\mathcal{W}$ acts non-trivially on $\mathfrak{a}$. This leads to the fact that $M=\operatorname{Fix}_{G}(p,F)$. ∎
Note that from these isomorphisms used in the previous proof it follows that any non-trivial coset of $\operatorname{Stab}_{G}(e,AK)/M$ acts non-trivially on the flat $AK$.

### v4

@@ 4.4 Flats
Let $G$ be a split Kac–Moody group over $\mathbb{K}$ of type $\mathbb{A}$ and equip the real form $\mathfrak{a}$ with its Euclidean symmetric space structure (cf. 2.2 1). Then the map
@@
Now, using that the action of $G$ on the symmetric space $G/K$ is given by an automorphism (cf. 4.4), we can assume that $g=e$. Let $X,Y\in\mathfrak{a}$ and use that for all $t\in A$ we have that $\Theta(t)=t^{-1}$ (cf. Section 3.5). Then one computes
@@
To check whether the image $\mu(gaK,ga^{\prime}K)$ is contained in $gAK$ for all $a,a^{\prime}\in A$, we do a straightforward calculation
$\mu(gaK,ga^{\prime}K)$ $=\tau(ga)\Theta(ga^{\prime})K$
$=\tau(ga)\Theta(ga^{\prime})K$
$=(ga)\Theta(ga)^{-1}\Theta(ga^{\prime})K$
$=ga\Theta(a)^{-1}\Theta(g)^{-1}\Theta(g)\Theta(a^{\prime})K$
$=gaaa^{\prime-1}K\in gAK.$
@@
In claim $5$ a reductive split real Lie group is used and the fact that it is a subgroup of $\operatorname{GL}_{m}(\mathbb{R})$. For $\mathbb{K}=\mathbb{C}$ we also have an algebraic group which is a subgroup of $\operatorname{GL}_{m}(\mathbb{C})$ (cf. [Bor91, Proposition 1.10]. As we have seen in the proof of 4.1, the involution $\Theta$ at the level of the matrix group can be understood as transpose, inversion, which is combined with complex conjugation (cf. [PR94, Theorem 3.7 and Proposition 3.9]). In the concrete calculation of the product $xy\cdot xy$, the upper left entry is equal to a sum of the squares of the absolute values of the entries in the first column. Now the arguments in claim $5$ work verbatim exactly as in the real case. ∎
@@ 4.4.1 Flats in the Group Model
In Section 4.3 we have defined two isomorphisms $\eta\colon G/K\to T(G/K)$ and $\zeta\colon T(G/K)\to\tau(G)$. If we now combine them, we obtain the following isomorphism $\hat{\tau}\colon G/K\to\tau(G)$, $\hat{\tau}(gK)=\tau(g)$ between the coset and the group model and that it is equivariant under the action of $G$. We are now interested in the image of a standard flat in the group model under this isomorphism:
@@
$=g\tau(A)\Theta(g)^{-1}.$
In the following we use the notation
$g\star\tau(A)\coloneqq g\tau(A)\Theta(g)^{-1}.$
Thus, the image of the standard flat $gAK$ in the group model is given by $g\star\tau(A)$. Therefore, denote by $F[g]\coloneqq g\star A\subseteq\tau(G)$ the standard flat of the group model (cf. [FHHK20, Proposition 5.4]). In order to establish a relationship between the flats in the symmetric space $G/K$, the twin apartments in the corresponding building $\Delta_{\pm}$ and the maximal tori of $G$, we need the following lemma.
@@
Let $G$ be a split Kac–Moody group over $\mathbb{K}$ of arbitrary type. Then
@@
Recall that in the case $\mathbb{K}=\mathbb{C}$ $T$ is connected and in the case $\mathbb{K}=\mathbb{R}$ $A$ is the identity component of $T$. This indicates that $N_{K}(T)<N_{K}(A)$, based on the fact that the conjugation is continuous.
For the reverse inclusion, consider an element normalizing $A$, then the conjugation with this element shifts $T$ to a maximal torus $T^{\prime}$ of $G$. But the standard maximal torus $T$ containing $A$ is unique (a sufficiently large subgroup of $T$ fixes a unique twin apartment which in turn determines $T$; cf. [Cap09, Lemma 4.9]). Therefore $T=T^{\prime}$, thus this element normalizes $T$. ∎
@@
Note that the strongly transitive action of $G$ on the twin building $\Delta_{\pm}$ implies the transitive $G$-action on the set of twin apartments; cf. [AB08, Lemma 6.73]. The second claim is thus an immediate consequence of the first.
Denote by $T_{max}$ the set of all maximal tori of $G$ and with $F_{std}$ the set of all standard flats in $G/K$. Then there is a $G$-equivariant map between these sets given by
@@
Together with the fact that $N_{G}(T)$ acts trivially by conjugation on $T_{max}$, we conclude that $G/N_{G}(T)$ can be identified with $T_{max}$ by mapping $gN_{G}(T)\mapsto gTg^{-1}$, in particular $f$ is surjective, see also the last point of 4.10.
@@
Note that $T$ is the stabilizer of the standard twin apartment and the group $G$ acts transitively on $T$ and the twin apartments. For the action of $G$ on $T$ the stabilizer of the standard torus is given by $N_{G}(T)$ and for the action of $G$ on the twin apartments is $N_{G}(T)$ the stabilizer of the standard twin apartment. Thus, for the twin apartments it can be deduced that the maximal tori of $G$ are exactly the chamberwise stabilizers of the twin apartments of $\Delta_{\pm}$. ∎
Corollary 4.13
Let $G$ be a split Kac–Moody group over $\mathbb{K}$ of arbitrary type. Then
$N_{G}(T)=\operatorname{Stab}_{G}(AK)\coloneqq\{g\in G\mid g(AK)=AK\}.$
Proof:
The inclusion $N_{G}(T)\subseteq\operatorname{Stab}_{G}(AK)$ follows from the proof of the theorem 4.12.
For the inverse inclusion, let an element $g$ of $\operatorname{Stab}_{G}(AK)$ be given, i.e. $g(AK)=AK$. Then calculate
$gTg^{-1}=g(AM)g^{-1}=AM=T,$
where we have used the decomposition of $T$ and the fact that $M\subseteq K$, see 3.14. ∎
@@ 5 Automorphism of Kac–Moody Symmetric Spaces
As a next step in developing the theory of symmetric spaces for Kac–Moody groups over $\mathbb{K}$ of type $\mathbb{A}$, one can analyze the automorphism group of the whole symmetric space and the local action on flats. For the non-affine real case, in chapter 6 of [FHHK20], the authors observe several statements for the global automorphism group on the symmetric space as well as for the local automorphism group with respect to the flats.
@@ 5.1 Global Action
We are going to use [FHHK20, Proposition 2.33]. This statement is shown for abstract symmetric spaces, hence we can apply it here. In order to use this statement, we will check the required conditions. In detail, it has to be checked that $\operatorname{Trans}(G/K)<G<\operatorname{Aut}(G/K)$, furthermore that $G$ acts transitively on maximal flats and that all maximal flats are Euclidean. The right inclusion of the first condition is obviously satisfied, and the left inclusion is shown in 4.7.
From 4.12 it follows that $G$ acts transitively on standard flats. Since all standard flats are maximal flats and all maximal flats according to 4.10 are Euclidean in our context, all assumptions of [FHHK20, Proposition 2.33] are fulfilled. ∎
A central assumption in [FHHK20] which allows to decompose an automorphism, is that the Kac–Moody group is centered (cf. 3.13), i.e. the Kac–Moody group is generated only by its root groups. To establish a similar results for non-centered Kac–Moody groups $G$ of type $\mathbb{A}$, we have to take care of the extended torus. In order to do this, we consider in the following section $G^{\dagger}\trianglelefteq G$.
@@
To do this, we need to recall some things from Section 3.2. First, the map $\eta$ from the Tits functor, further the fact, that an element of the root group $U_{\alpha}$ is given by $x_{\alpha}(r)$ for $r\in\mathbb{K}\in\{\mathbb{R},\mathbb{C}\}$ and $\alpha\in\Delta^{re}$. And lastly, that the torus of a Kac–Moody group $T=\mathfrak{T}_{\Lambda}(\mathbb{K})$ is isomorphic to $\operatorname{Hom}_{grp}(\Lambda,\mathbb{K}^{\times})$. Hence we can define the map
@@
From the discussion in [CM05, Section 8.2] follows that the sign automorphism is unique and from [Cap09, Section 8.2.1] follows that it is related to the Chevalley involution. This leads to the conclusion that for $\mathbb{K}=\mathbb{R}$ one has
@@
The case $\mathbb{K}=\mathbb{R}$ is proven in [FHHK20, Proposition 6.5], where it is formulated for any generalized Cartan-Matrix $\mathbb{A}$.
For the complex case, we recall the decomposition of an automorphism of $G^{\dagger}$ ([CM05, Theorem 2.4]): Denote with $c_{g}$ an inner automorphism w.r.t. an element $g\in G$, with $\omega$ the Chevalley involution, with $d$ a diagonal automorphism, with $\gamma$ a diagram automorphism and with $f$ a field automorphism. Then any automorphism $\alpha\in G^{\dagger}$ can be written as
@@
In the complex case, there is only one non-trivial continuous field automorphism $f$, the complex conjugation. By the definition of $\gamma$ and $d$ we know that $f$ commutes with them and we can deduce that
$\alpha=c_{g}\circ\omega\circ f\circ d\circ\gamma=c_{g}\circ\Theta\circ d\circ\gamma,$
where we have used that $\Theta$ is defined as $\omega\circ f=\Theta$.
Since the varioua subgroups generated by these automorphisms intersect trivially, it follows that the product of the generated subgroups is semi-direct. ∎
@@
  • Since we are interested in statements about Kac–Moody symmetric spaces, we also need to define a symmetric space with respect to the centered subgroup $G^{\dagger}$. In particular, let $K^{\dagger}\coloneqq G^{\dagger}\cap K$, and it follows directly $(G^{\dagger})^{\Theta}=K^{\dagger}$. Now define the corresponding Kac–Moody symmetric space
@@
We conclude this section by pointing out the following statement, which describes the automorphism group of a Kac–Moody symmetric space coming from the centered subgroup.
@@
Proof:
The proof is analogous to the proof for [FHHK20, Theorem 6.12 (i)]. Note that, by 5.4, the symmetric space $X^{\dagger}$ is the same symmetric space considered in [FHHK20]. Additionally, the underlying field and type of Kac–Moody groups are irrelevant as they utilize 5.3, a basic statement applicable to all cases. ∎
@@ 6 Local Action
  • Call a point $q\in F$ singular with respect to $p$, if there exists a second maximal flat $E\subseteq X$ such that $p,q\in E$ and $E\neq F$. If not, call $q$ regular. Further, let ${F^{\text{sing}}}(p)$ denote the subset of singular points of $(p,F)$ and ${F^{\text{reg}}}(p)$ the set of regular points. Hence the flat $F$ can be written as $F={F^{\text{reg}}}(p)\sqcup{F^{\text{sing}}}(p)$.
@@
From the non-degeneracy of B on $\mathfrak{h}$ we conclude that B is also non-degenerate on $\mathfrak{a}$. Therefore, the reflection $r_{\alpha}^{\vee}$ under $\rho_{KM}$ is the only B-orthogonal reflection on the hyperplane $H_{\alpha}$. In particular, this leads to a one-to-one correspondence between the real roots $\alpha\in\Delta^{re}$ and the hyperplanes $H_{\alpha}$.
@@
As we saw at the beginning of this section, we can also define singular points on $\mathfrak{a}$ by using the notion of singular points in the context of abstract symmetric spaces. To be able to compare the two terms, we need to make them compatible. Therefore, let $(e,AK)$ be the standard flat of the symmetric space $G/K$. Using the chart $\varphi_{e}$ (for the definition see 4.8), one can move the singular points of the flat $AK$ into the real form $\mathfrak{a}$, i.e.
@@
Since $G$ acts strongly transitively, see 5.1, any pointed maximal flat $(p,F)$ can be mapped to the standard pointed flat $(e,AK)$. Furthermore, the $G$ action preserves the decomposition into regular and singular points. Thus we observe an isomorphism
@@
In order to be able to use some corollaries from the appendix of [FHHK20], we need an additional assumption in the rest of this section: the size $n$ for the symmetrizable generalized Cartan matrix $\mathbb{A}\in\mathbb{Z}^{n\times n}$ must be at least $2$.
@@
Note that [FHHK20, Corollary A.37] is applicable here, since the arguments and statements that the authors used in the proof of this corollary are based on the fact that the bilinear form B coming from the generalized Cartan matrix $\mathbb{A}$ is non-degenerate and that $\rho_{KM}$ is faithful. All these properties are also satisfied in our context and thus we can apply the corollary.
@@
Therefore given a pointed flat $(p,F)$ and a local transformation $f\in\operatorname{GL}(p,F,{F^{\text{sing}}})$, then $f$ is a local automorphism if $\varphi^{-1}\circ f\circ\varphi$ is B-orthogonal for a chart $\varphi$ centered at $p$. We denote the subgroup of local automorphism by $\operatorname{Aut}(p,F)<\operatorname{GL}(p,F,{F^{\text{sing}}})$. By the previous discussion, we can decompose any local transformation into a product of a homothety and a B-orthogonal transformation. Thus we can decompose the group of local transformation as follows
@@
$\operatorname{Aut}(\Sigma)=\mathcal{W}\rtimes\operatorname{Aut}(\mathcal{W},S)$
@@ 6.1 Comparing the Weyl Groups
Note that $\operatorname{Fix}_{G}(p,U)$ is the same as $\operatorname{Fix}_{G}(U)$, but adding the $p$ in the notation makes it clear that we are considering an pointed flat here.
@@
Recall that the extended Weyl group $\widetilde{\mathcal{W}}\coloneqq\langle\tilde{s}_{i}\mid i\in I\rangle$ (cf. Section 3.2) surjects on $\mathcal{W}$.
@@
where $A^{\dagger}$ and $M^{\dagger}$ are defined as in 3.14. By the construction of $G^{\dagger}$ follows that $T^{\dagger}\cong(\mathbb{R}^{\times})^{n}$, where $n$ is the size of the Cartan matrix $\mathbb{A}$. Hence it follows by the construction of $M^{\dagger}$ (cf. 3.14) that $M^{\dagger}$ is the unique maximal finite subgroup of elements of order $2$. This implies the following
@@
Altogether we conclude that
@@
On the other hand, one calculates
$\nicefrac{{\widetilde{\mathcal{W}}M}}{{M}}\cong\nicefrac{{\widetilde{\mathcal{W}}}}{{(\widetilde{\mathcal{W}}\cap M)}}=\nicefrac{{\widetilde{\mathcal{W}}}}{{M^{\dagger}}}\cong\mathcal{W}.$
For the first isomorphism the isomorphism theorem is used, for the second the definition of $M^{\dagger}$ with the fact that $\widetilde{\mathcal{W}}\cap T=M^{\dagger}$ is used, and for the last isomorphism Equation 10 is used.
From this follows
@@
Now we analyze the stabilizer and fixator of $G/K$ with respect to the action of $G$ on $G/K$. Observe that the fixator of $e$ is given by $K$ and $N_{G}(T)$ is the stabilizer of $AK$ (see 4.13). Thus $\operatorname{Stab}_{G}(e,AK)=N_{K}(T)$ follows and by 6.4 follows that $\widetilde{\mathcal{W}}M=N_{K}(T)$.
Recall from 3.14 (A3) that $T=M\times A$, in particular $M$ centralizes $A$. Hence $M$ is a subgroup of $\operatorname{Fix}_{G}(e,AK)$. Using 6.4 and the isomorphism Equation 10 we conclude that
@@
So far we have established the isomorphism $\operatorname{Stab}_{G}(e,AK)/M\cong\mathcal{W}$. Since the representation $\rho_{KM}\colon\mathcal{W}\to\operatorname{GL}(\mathfrak{a})$ (cf. Equation 7) is faithful, each element different from the identity of $\mathcal{W}$ acts non-trivially on $\mathfrak{a}$. The isomorphism $\nicefrac{{\operatorname{Stab}_{G}(e,AK)}}{{M}}\cong\mathcal{W}$ from above hence implies that $M=\operatorname{Fix}_{G}(e,AK)$. ∎
Note that from these isomorphisms used in the previous proof it can be deduced that any non-trivial coset of $\operatorname{Stab}_{G}(e,AK)/M$ acts non-trivially on the flat $AK$.